OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^6/x^2. Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form.

OpenStudy (anonymous):

@RadEn @sachinrajsharma

OpenStudy (raden):

because the base is same, just subtract the power of both

OpenStudy (anonymous):

x^4 ?

OpenStudy (raden):

exactly, yep!

OpenStudy (anonymous):

x^6 ---- = x^(6-2) =x^4 x^2

OpenStudy (anonymous):

what is the expanded form?

OpenStudy (anonymous):

Part 1: x^6/x^2=x^(6-2)=x^4 When you divide monomials, you actually subtract their powers. Remember that division is related to subtraction.

OpenStudy (raden):

the expanding form mean : for x^6 = x * x * x * x * x * x for x^4 = x * x * x * x

OpenStudy (anonymous):

Show the expanded form and simplification of X^6/X^2 . Explain in your own words how you can simplify X^6/X^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification. I don't need to know how to simplify it i just need the answers of the questions... thanks oh and best answer gets 10 points....lol Answer Part 1: x^6/x^2=x^(6-2)=x^4 When you divide monomials, you actually subtract their powers. Remember that division is related to subtraction. Part 2: y^3/x^3y =y^2/x^3

OpenStudy (anonymous):

Did that help?

OpenStudy (raden):

so, without do cancel out the same terms we can simplify it by the property exponent, just subtract their powers.

OpenStudy (anonymous):

Thank you to both of you... I would like to give both of you medals .... i will give a medal to @RadEn and I will write thanks to @godgavemeyou (comment) and you will get a medal too